# Gradient Dynamics of Attention: How Cross-Entropy Sculpts Bayesian Manifolds

### v2

@@ Abstract.
Transformers empirically perform precise probabilistic reasoning in carefully constructed “Bayesian wind tunnels” and in large-scale language models, yet the mechanisms by which gradient-based learning creates the required internal geometry remain opaque. We provide a systematic first-order analysis of how cross-entropy training reshapes attention scores and value vectors in a transformer attention head. Our core result is an advantage-based routing law for attention scores,
@@
where $u_{i}$ is the upstream gradient at position $i$ and $\alpha_{ij}$ are attention weights. These equations induce a positive feedback loop in which routing and content specialize together: queries route more strongly to values that are above-average for their error signal, and those values are pulled toward the queries that use them. We show that this coupled specialization behaves like a two-timescale EM procedure: attention weights implement an E-step (soft responsibilities), while values implement an M-step (responsibility-weighted prototype updates), with queries and keys adjusting the hypothesis frame. Through controlled simulations, including a sticky Markov-chain task where we compare a closed-form EM-style update to standard SGD, we demonstrate that the same gradient dynamics that minimize cross-entropy also sculpt the low-dimensional manifolds identified in our companion work as implementing Bayesian inference. Further, in EM versus SGD comparison, EM converges much faster, and predictive entropy for both is close to minimum Bayesian entropy. This yields a unified picture in which optimization (gradient flow) gives rise to geometry (Bayesian manifolds), which in turn supports function (in-context probabilistic reasoning).
@@ 1. Introduction
Transformers have become the dominant architecture for sequence modeling, yet we still lack a mechanistic understanding of how gradient descent sculpts their internal representations. Recent work has shown that in controlled “Bayesian wind tunnels” a small transformer can reproduce analytic posteriors exactly, with keys forming orthogonal hypothesis axes, queries implementing progressive belief updates, and values unfurling along one-dimensional manifolds parameterized by posterior entropy (Agarwal et al., 2025). Our companion scaling paper extends this picture to production-scale models, showing similar geometric signatures in Pythia, Phi-2, and LLaMA.
These findings raise a natural question: Why does plain cross-entropy training produce the very geometric structures required for Bayesian inference? Understanding this requires moving from static geometry to gradient dynamics: how do attention scores, queries, keys, and values co-evolve during optimization?
This paper provides a mechanistic account. We analyze a single-head attention block trained with cross-entropy and derive all first-order gradients with respect to scores $s_{ij}$, queries $q_{i}$, keys $k_{j}$, and values $v_{j}$. We then interpret the resulting gradient flows and connect them to the Bayesian manifolds observed empirically.
@@ 1.1. Contributions
    where $b_{ij}=u_{i}^{\top}v_{j}$ is a compatibility term. Attention scores increase for positions whose compatibility exceeds the current attention-weighted mean, and decrease otherwise.
@@
    Toy experiments and EM vs. SGD comparison. In synthetic tasks, including a sticky Markov-chain sequence, we compare a closed-form EM-style update induced by our gradient analysis to standard SGD. EM reaches low loss, high accuracy, and sharp predictive entropy significantly faster; SGD converges to similar solutions but with slower and more diffuse routing. PCA visualizations of value trajectories reveal emergent low-dimensional manifolds.
@@
Clarification on “Bayesian inference.” Throughout this paper, “Bayesian inference” refers to the Bayesian posterior predictive over latent task variables—not a posterior over network weights. We show that cross-entropy training sculpts geometry that implements Bayesian filtering over in-context hypotheses, not Bayesian weight uncertainty.
@@ 3.4. Gradient with Respect to Scores
Under this definition, $A_{ij}>0$ indicates that increasing $s_{ij}$ (and hence $\alpha_{ij}$) would reduce the loss. Gradient descent therefore increases attention scores toward positions with positive advantage and decreases them otherwise.
This is an advantage-style gradient: scores increase for positions whose compatibility exceeds the current attention-weighted average, and decrease otherwise.
@@ 3.5. Gradients to $q_{i}$ and $k_{j}$
We next propagate through the score definition $s_{ij}=q_{i}^{\top}k_{j}/\sqrt{d_{k}}$.
For fixed $i$,
@@ 4. Coupled Dynamics and Specialization
We now unpack the implications of the gradient flows in Section Section 3, focusing on the interaction between routing (via scores and attention) and content (via values).
@@ 4.1. Advantage-Based Attention Reallocation
Equation (23) shows that for fixed query $i$, It is convenient to define an advantage quantity with the sign chosen to align with gradient descent:
$A_{ij}:=-\bigl(b_{ij}-\mathbb{E}_{\alpha_{i}}[b]\bigr).$
Under this definition, $A_{ij}>0$ indicates that increasing $s_{ij}$ (and hence $\alpha_{ij}$) would reduce the loss. Gradient descent therefore increases attention scores toward positions with positive advantage and decreases them otherwise.
Thus attention reallocates mass away from positions whose values are worse-than-average (i.e., $b_{ik}>\mathbb{E}_{\alpha_{i}}[b]$) for $u_{i}$ and toward those that are better-than-average, implementing an advantage-based routing rule.
@@ 4.2. Value Updates as Responsibility-Weighted Prototypes
so $v_{j}$ moves in the direction that jointly benefits all queries that attend to it, weighted by their attention.
In particular, if a small set of queries uses $j$ heavily, $v_{j}$ becomes a prototype that serves that subset. This is the core specialization mechanism: values adapt to the error landscape created by their users.
@@ 4.3. First-Order Effect on Contexts and Loss
The factor $\alpha_{ij}$ says: the more query $i$ relies on value $j$, the more $v_{j}$’s update affects $g_{i}$. The column weights $\{\alpha_{rj}\}$ aggregate contributions from all queries that share $j$. The inner products $u_{i}^{\top}u_{r}$ measure whether their error directions are aligned (helpful) or opposed (conflicting).
If many $u_{r}$ align with $u_{i}$, then $v_{j}$ can move in a direction that helps them all. If some are anti-aligned, $v_{j}$ must compromise, and specialization may become harder.
@@ 4.4. Feedback Loop and Specialization
    If $v_{j}$ is particularly helpful to $i$ (large negative $b_{ij}-\mathbb{E}_{\alpha_{i}}[b]$), then $s_{ij}$ increases and $\alpha_{ij}$ grows.
@@
    Larger $\alpha_{ij}$ gives $u_{i}$ more weight in $\bar{u}_{j}$, so $v_{j}$ moves further toward $u_{i}$.
@@
    As $v_{j}$ aligns with $u_{i}$, $b_{ij}=u_{i}^{\top}v_{j}$ increases, reinforcing step 1.
@@ 4.5. Geometric Illustration
Figure Figure 1 schematically illustrates the coupled dynamics: $v_{j}$ is pulled toward upstream gradient $u_{i}$, inducing a change in $g_{i}$ along a direction that reduces loss; the improved alignment feeds back into the score gradients.
@@ 5. EM-Like Two-Timescale Dynamics
The coupled dynamics derived above admit a useful analogy to Expectation–Maximization (EM), not as a literal optimization of an explicit latent-variable likelihood, but as a mechanistic correspondence between gradient flows and responsibility-weighted updates. Attention weights behave like soft responsibilities over latent sources, while value vectors act as prototypes updated under those responsibilities. Unlike classical EM, the updates here are driven by upstream gradients rather than observed data, and no standalone likelihood over values is optimized.
@@ 5.2. Values as Prototype Updates
with $u_{i}$ playing the role of residuals that need to be explained.
A key difference from classical mixture-model EM is the role played by $u_{i}$. In standard EM, centroids are updated using observed data points. Here, $u_{i}$ represents the residual error signal backpropagated from the loss. Values therefore move to better explain the error geometry induced by the current routing structure, rather than to maximize a likelihood over inputs. The analogy is structural rather than variational.
@@
  • Fast E-step: Score gradients quickly adjust $q_{i}$ and $k_{j}$ so that attention patterns allocate responsibility to helpful positions. In practice, attention often stabilizes early.
  • Slow M-step: Value vectors continue to move under residual error signals, refining the content manifold even after attention appears frozen. This continues until calibration and likelihood converge.
We can make this connection more explicit by considering a “manual EM” algorithm that alternates:
@@
In Section Section 7.2, we compare such an EM-like schedule to standard SGD on a sticky Markov-chain task and find that both converge to similar solutions, but the EM-like updates reach low loss and sharp, focused attention much faster.
@@ 5.4. Analogy Table
EM is an optimization procedure: it produces a point estimate $\theta^{\ast}$ maximizing (posterior) likelihood. A fully Bayesian treatment would instead integrate over $\theta$, but is intractable for transformers. Our analysis therefore lives at the EM/SGD level.
Crucially, however, our companion work (Agarwal et al., 2025) shows that the point-estimate parameters learned in this way support Bayesian computation in representation space: value manifolds, key frames, and query trajectories implement Bayesian belief updates in context. The present paper explains why cross-entropy and gradient descent naturally create these structures.
@@ 6.1. Value Manifold Unfurling
In wind-tunnel experiments (Agarwal et al., 2025), we observe that:
@@
  • Later in training, attention patterns appear stable, but value representations unfurl along a smooth curve; the first few principal components explain most variance, and the main axis correlates strongly with posterior entropy.
@@ 6.2. Hypothesis Frames and Key Orthogonality
If different subsets of queries consistently find different keys helpful, the corresponding gradient contributions push keys apart in $k$-space, encouraging approximate orthogonality of distinct hypothesis axes. Our wind-tunnel paper measures exactly this orthogonality and connects it to clean separation of competing hypotheses.
@@ 7.1. Toy Attention Simulation
  • We run manual update steps using the closed-form gradients from Section Section 3, with a fixed learning rate.
@@
    Attention heatmaps sharpen: mass concentrates on a few positions per query (Figures Figure 3, Figure 3).
@@
    Value vectors move coherently in a low-dimensional subspace; their trajectories in a PCA projection show emergent manifold structure (Figure Figure 5).
@@
    Cross-entropy loss decays smoothly (Figure Figure 5), with most gains occurring as specialization emerges.
@@
We generate sequences of length $T=2000$ over an 8-symbol vocabulary $\{0,\dots,7\}$ from a first-order Markov chain with self-transition probability $P(y_{t+1}=y_{t}\mid y_{t})=0.3$, otherwise transition to a different state with probability proportional to modulo distance of 7 with probability proportional to the distance from the previous state. With this circular distance, the nearby symbol gets higher probability than the distant one. Each symbol $y_{t}$ has an associated mean embedding in $\mathbb{R}^{20}$; the input at time $t$ is $x_{t}=\mu_{y_{t-1}}+\epsilon_{t}$ with $\epsilon_{t}\sim\mathcal{N}(0,I)$, so embeddings carry information about the previous symbol.
We train a single-head attention layer with $d_{x}=20$, $d_{k}=d_{q}=10$, $d_{v}=15$ for 1000 steps to predict $y_{t}$ from $(x_{1},\dots,x_{t})$.
@@
We compare two training schemes with causal masking:
@@
    Standard SGD. Vanilla gradient descent on all parameters with a fixed learning rate.
@@
    EM-like schedule. Alternating steps:
    • E-like step: forward pass to compute $\alpha_{ij}$ and $u_{i}$.
    • M-like step: update $v_{j}$ using the closed-form gradient $\Delta v_{j}=-\eta\sum_{i}\alpha_{ij}u_{i}$.
    • Small SGD step on $W_{Q},W_{K},W_{O}$ to adjust the latent assignment model.
@@ Results.
Table Table 1 summarizes the results after a fixed budget of steps.
Table 1. Sticky Markov-chain task: EM-like vs. standard SGD with stay_prob=0.3, after 1000 steps. EM achieves lower loss, higher accuracy, and sharper (lower entropy) predictive distributions. The theoretical Bayesian minimum predictive entropy is computed assuming $x_{t+1}=y(t)$. True theoretical predictive entropy would be slightly higher, since $x_{t}=y_{t-1}+\epsilon_{t}$, $\epsilon_{t}\sim\mathcal{N}_{d_{x}}(0,I)$, which increases entropy. The increase is small by the central limit theorem, since the noise $\mathcal{N}(0,I)$ is averaged over $d_{x}$ dimensions. The final row reports the KL divergence between EM and SGD predictive distributions.
Metric Value
final_loss_EM $1.8961$
final_loss_sgd $1.9177$
final_acc_EM $0.2690$
final_acc_sgd $0.2675$
final_entropy_EM $1.9076$
final_entropy_sgd $1.9276$
Theoretical Min entropy $1.8833$
final_KL(EM||SGD) $0.1019$
Figures Figure 7Figure 9 plot loss, accuracy, and entropy as a function of training step. As expected, EM and SGD methods converge to the same solution. This may not always happen due to local minima. EM converges much faster, with loss and entropy dropping sharply and accuracy rising quickly, while SGD progresses more slowly. This happens since $v_{j}$ are optimized directly in the function space, unlike SGD which optimizes in the parameter space. The argument is similar to functional gradient boost that converges faster than parameter updates (@(Friedman, 2001)). It is also interesting to note that even with this single layer and head architecture, the final empirical predictive entropy is very close to the theoretically minimum Bayesian entropy as described in Table 1. The theoretical entropy is higher due to addition of noise in defining $x_{t+1}$.
@@
To compare micro behavior of EM and sgd, we computed the movement of $v$ vector for EM and sgd. Since $T=2000$, we will have 2000 $v_{j}$ for each, EM and sgd. For clearer visualization we show a sample of 200 $v^{\prime}s$. The PCA projection plot of these 200 $v_{j}$ vectors trajectories is quite insightful. It shows (Figure Figure 9) that EM induces longer, more coherent trajectories from initial to final positions than SGD, consistent with stronger, more focused specialization. Red arrows (SGD) are shorter and more scattered; blue arrows (EM) are longer and more aligned, indicating that the EM-like updates exploit the analytic form of the value gradient more effectively.

### v3

@@ Abstract.
Our companion paper (Paper I) establishes that neural sequence models can implement exact Bayesian inference, with success depending on whether the architecture realizes the required inference primitives: belief accumulation, belief transport, and random-access binding. But how does gradient descent learn to implement these primitives? We provide a systematic first-order analysis of how cross-entropy training reshapes attention scores and value vectors. Our core result is an advantage-based routing gradient for attention scores,
@@
where $u_{i}$ is the upstream gradient at position $i$ and $\alpha_{ij}$ are attention weights. These equations induce a positive feedback loop: queries route more strongly to values that are above-average for their error signal, and those values are pulled toward the queries that use them. We show that this coupled specialization behaves like a two-timescale EM procedure: attention weights implement an E-step (soft responsibilities), while values implement an M-step (responsibility-weighted prototype updates). This EM-like dynamic is what enables the inference primitives: belief accumulation emerges from responsibility-weighted value updates; belief transport emerges from content-dependent routing that tracks evolving states; and random-access binding emerges from the query-key matching that allows retrieval by content. Through controlled simulations, we demonstrate that the same gradient dynamics that minimize cross-entropy also sculpt the low-dimensional Bayesian manifolds observed in Paper I. We further propose an abstract framework for content-based value routing that encompasses both attention and selective state-space models, conjecturing that advantage-based routing dynamics emerge in any architecture satisfying this framework—explaining why transformers and Mamba develop Bayesian geometry while LSTMs do not.
@@ 1. Introduction
Our companion paper (Paper I) establishes that neural sequence models can implement exact Bayesian inference—filtering and hypothesis elimination—in controlled “Bayesian wind tunnels.” The key finding is that success depends on whether the architecture realizes the required inference primitives: belief accumulation (integrating evidence), belief transport (propagating beliefs through dynamics), and random-access binding (retrieving hypotheses by content). Transformers realize all three primitives; Mamba realizes accumulation and transport; LSTMs realize only accumulation of static sufficient statistics; MLPs realize none. This taxonomy explains the empirical pattern: each architecture succeeds precisely on tasks demanding only the primitives it can implement. But how does gradient descent learn to implement these primitives? Why does cross-entropy training produce the geometric structures required for Bayesian inference—orthogonal key bases, progressive query alignment, low-dimensional value manifolds?
This paper as Lemma 2 (Mechanism). Paper I establishes existence: transformers can implement exact Bayesian inference, and different architectures realize different primitives. This paper establishes mechanism: when an architecture supports a given primitive, standard cross-entropy training induces gradient dynamics that reliably construct the corresponding probabilistic computation. Paper III establishes scaling: these mechanisms persist in production LLMs. We analyze a single-head attention block trained with cross-entropy and derive all first-order gradients with respect to scores $s_{ij}$, queries $q_{i}$, keys $k_{j}$, and values $v_{j}$. We show that the resulting gradient dynamics implement an implicit EM-like algorithm, and that this EM structure is precisely what enables the inference primitives to emerge.
@@ 1.1. Contributions
    where $b_{ij}=u_{i}^{\top}v_{j}$ is a compatibility term. Since gradient descent subtracts the gradient, scores decrease for positions whose compatibility exceeds the current attention-weighted mean (positive gradient), and increase for positions below average (negative gradient).
@@
    Toy experiments and EM vs. SGD comparison. In synthetic tasks, including a sticky Markov-chain sequence, we compare an EM-motivated learning rate schedule (with larger LR for values) to standard SGD. The EM-like schedule reaches low loss, high accuracy, and sharp predictive entropy significantly faster; SGD converges to similar solutions but with slower and more diffuse routing. PCA visualizations of value trajectories reveal emergent low-dimensional manifolds.
@@
Clarification on “Bayesian inference.” Throughout this paper, “Bayesian inference” refers to the Bayesian posterior predictive over latent task variables—not a posterior over network weights. Specifically, we address filtering and hypothesis elimination in tasks whose posteriors factorize sequentially, not general Bayesian model selection. We show that cross-entropy training sculpts geometry that implements these computations over in-context hypotheses.
@@ 3.4. Gradient with Respect to Scores
Under this definition, $A_{ij}>0$ indicates that increasing $s_{ij}$ (and hence $\alpha_{ij}$) would reduce the loss. Gradient descent therefore increases attention scores toward positions with positive advantage and decreases them otherwise. This is an advantage-style gradient: since gradient descent subtracts the gradient, scores decrease for positions whose compatibility $b_{ik}$ exceeds the attention-weighted average (positive gradient), and increase for positions below average (negative gradient). Since high $b_{ik}=u_{i}^{\top}v_{k}$ indicates $v_{k}$ aligns with the error direction, this reallocates attention away from harmful values toward helpful ones.
@@ 3.5. Gradients to $q_{i}$ and $k_{j}$
We next propagate through the score definition $s_{ij}=q_{i}^{\top}k_{j}/\sqrt{d_{k}}$. For fixed $i$,
@@ 4. Coupled Dynamics and Specialization
We now unpack the implications of the gradient flows in Section 3, focusing on the interaction between routing (via scores and attention) and content (via values).
@@ 4.1. Advantage-Based Attention Reallocation
Equation (23) shows that for fixed query $i$, attention reallocates mass away from positions whose values are worse-than-average (i.e., $b_{ik}>\mathbb{E}_{\alpha_{i}}[b]$) for $u_{i}$ and toward those that are better-than-average. Recall the advantage $A_{ij}:=-(b_{ij}-\mathbb{E}_{\alpha_{i}}[b])$ defined earlier: gradient descent increases attention scores toward positions with positive advantage, implementing an advantage-based routing rule.
@@ 4.2. Value Updates as Responsibility-Weighted Prototypes
so $v_{j}$ moves in the direction that jointly benefits all queries that attend to it, weighted by their attention. In particular, if a small set of queries uses $j$ heavily, $v_{j}$ becomes a prototype that serves that subset. This is the core specialization mechanism: values adapt to the error landscape created by their users.
@@ 4.3. First-Order Effect on Contexts and Loss
The factor $\alpha_{ij}$ says: the more query $i$ relies on value $j$, the more $v_{j}$’s update affects $g_{i}$. The column weights $\{\alpha_{rj}\}$ aggregate contributions from all queries that share $j$. The inner products $u_{i}^{\top}u_{r}$ measure whether their error directions are aligned (helpful) or opposed (conflicting). If many $u_{r}$ align with $u_{i}$, then $v_{j}$ can move in a direction that helps them all. If some are anti-aligned, $v_{j}$ must compromise, and specialization may become harder.
@@ 4.4. Feedback Loop and Specialization
    If $v_{j}$ is particularly helpful to $i$ (large negative $b_{ij}-\mathbb{E}_{\alpha_{i}}[b]$, meaning $v_{j}$ points away from the error direction $u_{i}$), then $s_{ij}$ increases and $\alpha_{ij}$ grows.
@@
    Larger $\alpha_{ij}$ gives $u_{i}$ more weight in $\bar{u}_{j}$, so $v_{j}$ moves further away from $u_{i}$ (since $\Delta v_{j}=-\eta\bar{u}_{j}$).
@@
    As $v_{j}$ moves away from $u_{i}$, $b_{ij}=u_{i}^{\top}v_{j}$ decreases further, increasing $A_{ij}$ and reinforcing step 1.
@@ 4.5. Geometric Illustration
Figure 1 schematically illustrates the coupled dynamics: $v_{j}$ is pulled toward upstream gradient $u_{i}$, inducing a change in $g_{i}$ along a direction that reduces loss; the improved alignment feeds back into the score gradients.
@@ 5. EM-Like Two-Timescale Dynamics
The coupled dynamics derived above admit a useful analogy to Expectation–Maximization (EM), not as a literal optimization of an explicit latent-variable likelihood, but as a mechanistic correspondence between gradient flows and responsibility-weighted updates. Attention weights behave like soft responsibilities over latent sources, while value vectors act as prototypes updated under those responsibilities. Unlike classical EM, the updates here are driven by upstream gradients rather than observed data, and no standalone likelihood over values is optimized. We emphasize that this is an interpretive framework: we do not derive a surrogate objective with guaranteed monotonic improvement. The value lies in the geometric intuition it provides for understanding specialization dynamics.
@@ 5.2. Values as Prototype Updates
with $u_{i}$ playing the role of residuals that need to be explained. A key difference from classical mixture-model EM is the role played by $u_{i}$. In standard EM, centroids are updated using observed data points. Here, $u_{i}$ represents the residual error signal backpropagated from the loss. Values therefore move to better explain the error geometry induced by the current routing structure, rather than to maximize a likelihood over inputs. The analogy is structural rather than variational.
@@
  • E-step (routing): Score gradients adjust $q_{i}$ and $k_{j}$ so that attention patterns allocate responsibility to helpful positions. In practice, attention often stabilizes relatively early.
  • M-step (values): Value vectors continue to move under residual error signals, refining the content manifold even after attention appears frozen. This continues until calibration and likelihood converge.
Both updates occur simultaneously from each gradient computation. The “two-timescale” behavior emerges from relative gradient magnitudes and coupling structure, not from explicit alternation. We can make this connection more explicit by considering a “manual EM” algorithm that alternates:
@@
In Section 7.2, we compare such an EM-like schedule to standard SGD on a sticky Markov-chain task and find that both converge to similar solutions, but the EM-like updates reach low loss and sharp, focused attention much faster.
@@ 5.4. Analogy Table
EM is an optimization procedure: it produces a point estimate $\theta^{\ast}$ maximizing (posterior) likelihood. A fully Bayesian treatment would instead integrate over $\theta$, but is intractable for transformers. Our analysis therefore lives at the EM/SGD level. However, our companion work (Agarwal et al., 2025) shows that the point-estimate parameters learned in this way support Bayesian computation in representation space: value manifolds, key frames, and query trajectories implement Bayesian belief updates in context. The present paper explains why cross-entropy and gradient descent naturally create these structures.
@@ 6.1. Value Manifold Unfurling
In wind-tunnel experiments (Agarwal et al., 2025), which directly measure the geometric predictions made here:
@@
  • Later in training, attention patterns appear stable, but value representations unfurl along a smooth curve; PC1 explains 84–90% of variance, strongly correlated with posterior entropy ($|r|>0.9$).
@@ 6.2. Hypothesis Frames and Key Orthogonality
If different subsets of queries consistently find different keys helpful, the corresponding gradient contributions push keys apart in $k$-space, encouraging approximate orthogonality of distinct hypothesis axes. Our wind-tunnel paper measures exactly this orthogonality: mean off-diagonal cosine similarity drops to 0.05–0.06 versus 0.08 for random vectors, a statistically significant reduction across multiple seeds ($p<0.001$).
@@ 7.1. Toy Attention Simulation
  • We run manual update steps using the closed-form gradients from Section 3, with a fixed learning rate.
@@
    Attention heatmaps sharpen: mass concentrates on a few positions per query (Figure 3, Figure 3).
@@
    Value vectors move coherently in a low-dimensional subspace; their trajectories in a PCA projection show emergent manifold structure (Figure 5).
@@
    Cross-entropy loss decays smoothly (Figure 5), with most gains occurring as specialization emerges.
@@
We generate sequences of length $T=2000$ over an 8-symbol vocabulary $\{0,\dots,7\}$ from a first-order Markov chain with self-transition probability $P(y_{t+1}=y_{t}\mid y_{t})=0.3$, otherwise transition to a different state with probability proportional to modulo distance of 7 with probability proportional to the distance from the previous state. With this circular distance, the nearby symbol gets higher probability than the distant one. Each symbol $y_{t}$ has an associated mean embedding in $\mathbb{R}^{20}$; the input at time $t$ is $x_{t}=\mu_{y_{t-1}}+\epsilon_{t}$ with $\epsilon_{t}\sim\mathcal{N}(0,I)$, so embeddings carry information about the previous symbol. We train a single-head attention layer with $d_{x}=20$, $d_{k}=d_{q}=10$, $d_{v}=15$ for 1000 steps to predict $y_{t}$ from $(x_{1},\dots,x_{t})$.
@@
We compare two training schemes with causal masking and matched total parameter updates per step:
@@
    Standard SGD. Vanilla gradient descent on all parameters with learning rate $\eta=0.01$.
@@
    EM-like schedule. Parameter-specific learning rates: $\eta_{V}=0.1$ for value parameters, $\eta_{\text{routing}}=0.01$ for $W_{Q},W_{K},W_{O}$. All updates are computed from a single forward-backward pass and applied simultaneously.
The EM-like schedule uses a larger learning rate for values ($10\times$) to accelerate value specialization, compensating for the fact that value gradients are typically smaller in magnitude than routing gradients early in training. Both methods perform one gradient computation per step, with updates applied simultaneously (not alternating) using the respective learning rates.
@@ Results.
Table 1 summarizes results across 5 random seeds after 1000 training steps.
Table 1. Sticky Markov-chain task: EM-like vs. standard SGD (mean $\pm$ std over 5 seeds). EM achieves significantly lower loss (paired $t$-test, $p=0.003$), higher accuracy, and sharper predictive distributions. Critically, EM reaches SGD’s final loss level in only $430\pm 143$ steps—a $2.3\times$ convergence speedup. The theoretical Bayesian minimum entropy is computed from the transition matrix; actual minimum is slightly higher due to input noise.
Metric EM-like SGD
Final loss $1.970\pm 0.034$ $2.058\pm 0.007$
Final entropy $1.998\pm 0.035$ $2.077\pm 0.001$
Final accuracy $0.244\pm 0.016$ $0.200\pm 0.009$
Steps to SGD level $430\pm 143$ 1000
Theoretical Min Entropy 1.829
Figure 7Figure 9 plot loss, accuracy, and entropy as a function of training step. EM-like training converges significantly faster than SGD: it reaches SGD’s final loss in under half the steps ($2.3\times$ speedup), with loss and entropy dropping sharply while accuracy rises quickly. This acceleration occurs because the EM-motivated learning rate schedule (with larger LR for values) allows values to specialize more rapidly, whereas uniform SGD must balance routing and content updates at the same rate. Even with this single-layer, single-head architecture, both methods approach the theoretical Bayesian minimum (1.829 bits), with EM converging closer (1.998 vs 2.077 bits entropy). The residual gap of $\sim$0.17 bits from the theoretical minimum likely reflects the limited capacity of a single-head architecture to fully capture the transition structure; concurrent work (Agarwal et al., 2025) shows that deeper architectures close this gap to <0.01<0.01 bits on similar tasks.
@@
To compare micro behavior of EM and SGD, we computed the movement of $v$ vectors for EM and SGD. Since $T=2000$, we will have 2000 $v_{j}$ for each method. For clearer visualization we show a sample of 200 $v$’s. The PCA projection plot of these 200 $v_{j}$ vector trajectories is quite insightful. It shows (Figure 9) that EM induces longer, more coherent trajectories from initial to final positions than SGD, consistent with stronger, more focused specialization. Red arrows (SGD) are shorter and more scattered; blue arrows (EM) are longer and more aligned, indicating that the EM-like updates exploit the analytic form of the value gradient more effectively.